1- find the next four terms of an arithmetic sequence 9,16,23........
2-determine the first three terms of arithmetic progression of which the 5th term is (7+9x) and d=2x
1
Expert's answer
2019-11-25T12:33:58-0500

1.

"a_1=9,\;\; a_2=16,\;\;a_3=23,\;\;d=16-9=7."

"a_4=23+7=30,"

"a_5=30+7=37,\\
a_6=37+7=44,\\
a_7=44+7=51."


2.

"a_5=a_1+4d,\\
a_1=a_5-4d=7+9x-2x*4=7+x,\\
a_2=a_1+d=7+x+2x=7+3x,\\
a_3=a_2+d=7+3x+2x=7+5x."